Some References:

M. Davis, T. Januszkiewicz, R. Scott, Nonpositive curvature of=20 blow-ups. Selecta Math. 4 (1998) 491-547

S. Devadoss, Tessellations of moduli spaces and the mosaic=20 operad, in Homotopy invariant algebraic structures, Contemp. Math. =

239 (1998) 91-114

K. Fukaya, Y-G Oh, H. Ohta, K. Ono, Lagrangian=20 intersection Floer theory: anomaly and obstruction, Kyoto, Dept. of = Math.=20 00-17

E. Getzler, Operads and moduli spaces of genus 0 Riemann=20 surfaces, in The moduli space of curves, 199–230, Progr. Math. = 129=20 (1995) 199-230, Birkhäuser

——, M. Kapranov, Cyclic operads and cyclic homology, in = Geometry,=20 Topology, and Physics for Raoul Bott (1994)

C. Kassel, Quantum groups, Springer Graduate Texts 155 (1995)

S. Keel, Intersection theory of moduli space of stable=20 N-pointed curves of genus zero. Trans. AMS 330 (1992) 545-574

P. Lochak, L. Schneps, The Grothendieck-Teichmüller group and automorphisms of braid groups, in The Grothendieck theory of dessins d’enfants, London Math. Soc. Lecture Notes 200 (1994) = 323-358=20

J. Moser, Various aspects of integrable Hamiltonian systems, in Dynamical systems (Bressanone, 1978) 233–289, Progr. Math. 8,=20 Birkhäuser (1980)

T.H. Parker, J.G. Wolfson, Pseudo-holomorphic maps and=20 bubble trees. J. Geom. Anal. 3 (1993) 63-98

M. Yoshida, The democratic compactification of configuration spaces of point sets on the real projective line, Kyushu J. of Math. 50 = (1996)=20 493-512